What are the present values for the 3 options of payouts??Congratulations! You have won a state lottery. The state lottery offers you the following (after-tax) payout options: BE: (Cl
Answer #1

Calculate present value

Option #1 = 14000000*0.56743 = $7944020

Option #2 = 2150000*3.60478 = $7750277

Option#3 = 13000000*0.71178 = 9253140

So Option #3 would be opt